I have one more that I am struggling with, it is a word problem. A farmer has 100 acres which he can plant corn or oats. Each acre of oats requires $18 capital and 6hrs. of labor. Labor is $8/hr. The farmer has $2100 for capital and $2400 for labor. If the revenue is $55 from each acre of oats and $125 from each acre of corn, what planting combinaton will produce the greatest total profit? List combination and the profit.
Answer by stanbon(57967)

A farmer has 100 acres which he can plant corn or oats.
Each acre of oats requires $18 capital and 6hrs. of labor.
Labor is $8/hr.
The farmer has $2100 for capital and $2400 for labor.
If the revenue is $55 from each acre of oats and $125 from each acre of corn, what planting combinaton will produce the greatest total profit?
List combination and the profit.
-----
Acre: oats + corn <= 100
Revenue = 55o+125c
capital = (18+6*8)o + ?c <= 2100
labor = (6*18)o + ?c <= 2400
---------------------
Note: You have not listed capital and labor costs for corn.
----
Note: Profit = Revenue - Cost
Find the max for the profit equation.
============================================
Cheers,
Stan H.